# DRIFT: Decompose, Retrieve, Illustrate, then Formalize Theorems

### v1

@@ 1 Introduction
Autoformalization is formulated as a translation task that aims to translate natural language mathematical descriptions into machine-verifiable statements written in formal languages, such as Coq (Barras et al., 1997), Isabelle (Paulson, 1994), and Lean (De Moura et al., 2015). Previous work has shown that accurate autoformalization is a critical step towards developing automated theorem proving systems (Lin et al., 2025b; Chen et al., 2025; Xin et al., 2024; Lin et al., 2025c), and ultimately assist mathematicians in new discoveries (Gouëzel & Shchur, 2019; Leang et al., 2025). Despite recent progress of Large Language Models (LLMs) in informal mathematical reasoning (Ahn et al., 2024; Setlur et al., 2024; Luong & Lockhart, 2025), formal reasoning presents distinct challenges. The strict syntax and necessity for precise alignment between informal concepts and formal definitions mean that even powerful pretrained models often fail at autoformalization tasks off-the-shelf (Wu et al., 2025).
@@
Although synthetic data generation could enable the fine-tuning of LLMs for autoformalization (Jiang et al., 2023; Lin et al., 2025a; Wang et al., 2024; Ying et al., 2024), the knowledge cut-off issue raised by updating formal libraries like Mathlib (Mathlib Community, 2020) makes finetuned models prone to hallucinating non-existent formal objects that have been renamed, reorganized or deprecated (Baanen et al., 2025). Early retrieval-augmented methods addressed this by retrieving similar theorems from external libraries to provide useful syntactic structure and compositional examples. However, their practical utility as exemplars has been limited by the retrieval methods used to find them. These methods often lack task-specific training data and rely on general-purpose techniques like keyword searching (Agrawal et al., 2022), k-NN (Azerbayev et al., 2023), or pretrained dense retrievers (Zhang et al., 2024). Liu et al. (2025) introduced dependency retrieval task and developed the RAutoformalizer (RAuto) framework. Similar to the premise selection for proof generation (Yang et al., 2023), this paradigm enables the training of specialized retrievers to identify the exact definitions that formal statements require. However, this new approach created a key trade-off: focusing on individual components meant losing the valuable context provided by full theorem statements. Based on this, we identify two main limitations in the current approach to retrieval-augmented autoformalization: Informal statements are often dense and multifaceted. This underlying complexity of queries makes them suboptimal as direct queries for retrieving the precise, atomic definitions required for formalization. Additionally, even when the correct formal definitions are retrieved, models often lack the knowledge of contextual usage required to correctly structure and integrate them into the formal statement.
@@
In information retrieval, query enhancement techniques like query expansion (Chan et al., 2024), pseudo-document generation (Gao et al., 2023), and neural query rewriting (Wang et al., 2025) have demonstrated effectiveness of reformulating queries to provide more semantic information. Query decomposition has proven particularly useful for multi-hop question answering, as it matches the granularity of the dense query statements with the indexed documents (Ammann et al., 2025). In addition to retrieving correct premises, providing rich context like exemplar proofs can guide proof generation effectively (He et al., 2024; Thakur et al., 2024; Thompson et al., 2025). Despite advances in adjacent domains, these techniques have not yet been systematically applied to dependency retrieval for autoformalization, which still relies on monolithic queries and provides context-free definitions.
@@
We propose Drift, a novel framework depicted in Figure 1, that enhances retrieval-augmented autoformalization by adapting query decomposition and context augmentation to address unique challenges of theorem autoformalization. To address complex queries, we adapt query decomposition to break down complex informal statements into a series of simpler, atomic sub-queries. Each sub-query targets a single mathematical concept, transforming a multifaceted retrieval problem into focused, precise searches. Second, to illustrate the usage of the dependencies, we contextualize the retrieved definitions with demonstrative theorems, giving the model concrete examples of syntax and application patterns.
@@
Contributions. 1) We introduce Drift 111The code and models are available at https://github.com/huawei-noah/DRIFT. (Decompose, Retrieve, Illustrate, then Formalize Theorems), a decomposition-driven retrieval augmented formalization framework that breaks down informal mathematical statements into atomic sub-queries and contextualizes retrieved premises with demonstrative theorems, transforming monolithic retrieval into a process aligned with the dependency structure of formal mathematics. 2) Our experiments establish new state-of-the-art in dependency retrieval and autoformalization on ProofNet and ConNF across models, with exceptional performance on out-of-distribution ConNF benchmark. 3) Through systematic analysis, we establish that the utility of retrieved dependencies is conditioned on the gap between a model’s parametric knowledge and the statement complexity. These insights reveal critical design considerations for retrieval-augmented systems and point toward the necessity of adaptive strategies that can assess when external knowledge genuinely complements model capabilities.
@@ 2 Related Work
Retrieval-augmented Autoformalization. Early retrieval-augmented autoformalization methods retrieved similar theorems as few-shot examples: ProofNet (Azerbayev et al., 2023) employs k-NN search, while MS-RAG (Zhang et al., 2024) uses informal-to-informal retrieval with iterative refinement. LTRAG (Hu et al., 2025) retrieves thought-guided theorem pairs for neuro-symbolic formalization. Liu et al. (2025) established this paradigm by introducing “dependency retrieval”, a premise selection task specialized for autoformalization: given an informal statement, retrieve the precise set of formal objects and definitions that are required for its autoformalization from a library (e.g. Mathlib). While RAutoformalizer (RAuto) (Liu et al., 2025) demonstrated improvements over non-retrieval methods, evaluation revealed a significant gap compared to oracle systems with ground-truth dependencies. CRANF (Lu et al., 2025) attempts conceptual mapping between abstraction levels. Critically, all existing approaches treat complex statements as monolithic queries, failing to identify distinct mathematical concepts within them.
Query Decomposition and Enhancement. Query enhancement has proven effective across retrieval tasks. Query2Doc (Wang et al., 2023) and HyDE (Gao et al., 2023) generate pseudo-documents to expand semantic coverage. SemanticSearch (Gao et al., 2024) augment the informal statement by prompting an LLM to translate the query into a detailed statement with both informal and formal expressions. However, they only evaluated with similar theorem retrieval but not on the downstream formalization. More relevant to our work, query decomposition has shown success in multi-hop question answering (Ammann et al., 2025), where breaking complex questions into sub-queries improves retrieval of distinct information aspects. Zhao et al. (2023) and Jiang et al. (2022) applied similar decompositions to theorem proving, showing the effectiveness of divide and conquer in formal math. Despite these successes, no prior work has applied decomposition to informal mathematical statements for autoformalization.
@@ 3 Methodology
We introduce Drift (Decompose, Retrieve, Illustrate, then Formalize Theorems), a novel four-stage method designed to address the two main limitations of previous retrieval-augmented formalization methods, 1) handling the complexity of informal statements and 2) the lack of demonstrative examples for retrieved formal objects. As a first step, an LLM decomposes the informal statement into a set of smaller, granular sub-queries (§3.1). These queries then guide the retrieval of dependent premises from a formal library such as Mathlib (§3.2). In a subsequent, bottom-up illustration step, we find theorems that utilize the retrieved premises, providing in-context demonstrations of their application (§3.3). Finally, the LLM formalizes the original statement, conditioned on all retrieved premises and theorems (§3.4). This process is visualized in Figure 1.
@@ 3.1 Decompose
Standard retrieval methods often treat complex informal statements as monolithic queries and simply embed the entire statement. This approach disregards the rich semantic structure within a statement, which may contain multiple distinct mathematical concepts. Their complexity means that a statement could be under-specified, ambiguous and/or simply too information-dense to be used as is. Compressing the entire statement’s meaning into a single dense vector creates a representational bottleneck, risking the loss of nuanced details and focusing the retrieval on only the most salient concepts. We hypothesize that by decomposing an informal statement into its constituent concepts, we can perform a more granular and accurate retrieval for each concept.
@@
Each sub-query is designed to isolate a single mathematical concept. As illustrated in Figure 1, the component is a natural language phrase describing the concept (e.g., “A prime number of the form is a prime that leaves remainder 3 when divided by 4”) while is a predicted formal representation for that concept (e.g., “ : with the conditions Nat.Prime and % , where % denotes the modulo operation on natural numbers.”). Appending this predicted formal name serves a dual purpose. First, it probes the LLM’s parametric knowledge, providing a syntactic “anchor” for the concept. Second, it allows the retriever to jointly leverage the semantics of the natural language phrase and the syntactic cues from to identify the correct premise even in the presence of minor inaccuracies in the predicted formal representations.
@@ 3.2 Retrieve
The Retrieve module is designed to identify dependent premises from the library that correspond to the concepts isolated in each sub-query. This one-to-one mapping is visualized in the ② panel of Figure 1, which shows each sub-query () being linked to a formal object (e.g., PNat.Prime). To accomplish this, we implement a dense passage retriever (Karpukhin et al., 2020) using a BGE-M3 (Chen et al., 2024) encoder (), finetuned on the dependency prediction task as introduced by Liu et al. (2025). This training objective aligns the formal dependencies with their informal statement, thereby making semantic similarity a strong proxy for logical dependency. We achieve this by encoding queries and formal library objects into a shared -dimensional embedding space. The vector representations for all formal objects are pre-computed in an offline step and stored in an efficient search index. At inference, each sub-query is encoded into a vector and the closest dependent premise is identified by finding the library object that maximizes the cosine similarity score, with the query vector. This is formally defined as:
@@ 3.3 Illustrate
Retrieving useful formal definitions is a necessary first step, however, it is not sufficient for successful autoformalization. For instance, a retrieved definition like def ZMod : → Type tells the model what the concept is, but provides no further guidance on its practical application such as the syntax for declaring a variable of that type (a : ZMod p). This gap between definition and usage is a primary source of syntactic and structural errors in LLM-generated formal statements.
The Illustrate module is designed to bridge this gap by providing examples of formal object usage, visualized in Figure 1 (panel ③). Given a premise like ZMod from the “Retrieve” step, the module selects illustrative theorems, e.g. “Theorem 2” which demonstrates the correct application of ZMod. The module takes the set of retrieved premises , and a budget , as input, and outputs a small set of illustrative theorems , where . The selection process is a greedy algorithm designed to maximize the coverage of the input premises as follows. First, we compile a candidate set of all theorems in library that utilize at least one of the retrieved premises .
In order to ensure relevance and provide a deterministic tie-breaker, we pre-sort this candidate set by the semantic similarity of each theorem’s informal statement to the original informal statement . The cosine similarity is computed using vectors from the same encoder as the retrieval stage: .
@@ 4.1 Benchmarks
We evaluate Drift on three distinct autoformalization benchmarks to test its in-distribution, self-contained and out-of-distribution performance. While the experiments are conducted in Lean, our framework is language-agnostic and adaptable to other formal systems with structured libraries.
@@
MiniF2F (Self-Contained). We use the MiniF2F-test set (Zheng et al., 2021) to evaluate the framework on self-contained problems. This benchmark consists of 224 Olympiad-style theorems with a notably low average of just 0.43 dependencies from Mathlib.222We removed duplicate and low-quality pairs that do not compile with our Lean version (4.18.0). MiniF2F-test serves as a boundary condition, testing the model’s core formalization capabilities and its robustness against potentially distracting context when retrieval is not strictly necessary.
ConNF (Out-of-Distribution). The OOD challenge refers to evaluation scenarios where neither the retrieval model nor the formalization model has been exposed to the formal objects used in the test data.333The knowledge cutoff dates of the models we used are June 2024, March 2025 and July 2025, for GPT-4.1, Claude-Opus-4, and DeepSeek-3.1, respectively. Though there is a risk of exposure to the formalizer which cannot be controlled or confirmed, from our zero-shot results on ConNF, it is hypothesised that the models have not been extensively trained on this benchmark. We therefore evaluate on ConNF, a benchmark curated by Liu et al. (2025) through topological informalization to test OOD generalization. This benchmark is based on the con-nf library and is not integrated with Mathlib.444The con-nf library is available at https://github.com/leanprover-community/con-nf. It formalizes a consistency proof for Quine’s New Foundations (Quine, 1951), established by Holmes & Wilshaw (2015) and contains 961 research-level theorems requiring an average of 3.92 premises from its distinct library of 1,348 formal objects. ConNF provides a rigorous test of Drift’s ability to generalize to a novel mathematical domain (Zhang et al., 2024).
@@ 4.2 Baselines
We evaluate our proposed framework against three key baselines representing zero-shot, state-of-the-art retrieval, and oracle-level performance. We note that concurrent work, CRAMF (Lu et al., 2025), was not publicly available for comparison at the time of our experiments. The no retrieval (zero-shot) baseline establishes a performance floor. The autoformalization model receives only the informal statement as input, without access to any retrieved context. DPR (RAuto) represents the current state-of-the-art. We compare our method to the dependency retrieval module from the RAutoformalizer framework (Liu et al., 2025), which is a finetuned dense passage retriever (DPR) (Karpukhin et al., 2020).555Model available at https://huggingface.co/purewhite42/dependency_retriever_f. The top-5 retrieved premises are provided to the formalizer model as augmented context. The oracle* (retrieval) setting provides an approximate upper bound for retrieval. The model is provided with ground-truth dependencies () for each problem, simulating a ’perfect’ retriever, as defined by Liu et al. (2025). We mark this setting with an asterisk (*) to denote that this oracle is in fact imperfect because the provided dependencies are not necessarily optimal or exhaustive for formalization and do not necessarily lead to the best autoformalization performance as we discuss in Section 5, i.e. some of our settings are actually outperforming this imperfect oracle.

### v3

@@ 1 Introduction
Autoformalization is formulated as a translation task that aims to translate natural language mathematical descriptions into machine-verifiable statements written in formal languages, such as Coq (Barras et al., 1997), Isabelle (Paulson, 1994), and Lean (De Moura et al., 2015). Previous work has shown that accurate autoformalization is a critical step towards developing automated theorem proving systems (Lin et al., 2025b; Chen et al., 2025; Xin et al., 2024; Lin et al., 2025c), and ultimately assisting mathematicians in new discoveries (Gouëzel & Shchur, 2019; Leang et al., 2025). Despite recent progress of Large Language Models (LLMs) in informal mathematical reasoning (Ahn et al., 2024; Setlur et al., 2024; Luong & Lockhart, 2025), formal reasoning presents distinct challenges. The strict syntax and necessity for precise alignment between informal concepts and formal definitions mean that even frontier LLMs often fail at autoformalization tasks off-the-shelf (Wu et al., 2025).
@@
Although synthetic data generation could enable the finetuning of LLMs for autoformalization (Jiang et al., 2023; Lin et al., 2025a; Wang et al., 2024; Ying et al., 2024), the knowledge cut-off issue raised by updating formal libraries like Mathlib (Mathlib Community, 2020) makes finetuned models prone to hallucinating non-existent formal objects that have been renamed, reorganized, or deprecated (Baanen et al., 2025). Early retrieval-augmented methods addressed this by retrieving similar theorems from external libraries to provide useful syntactic structure and compositional examples. However, their practical utility as exemplars has been limited by the retrieval methods used to find them. These methods often lack task-specific training data and rely on general-purpose techniques like keyword searching (Agrawal et al., 2022), k-NN (Azerbayev et al., 2023), or pretrained dense retrievers (Zhang et al., 2024). An advance is the introduction of the “dependency retrieval” task by Liu et al. (2025) with the RAutoformalizer (RAuto) framework. Similar to the premise selection for proof generation (Yang et al., 2023), this paradigm enables the training of specialized retrievers to identify the exact definitions that formal statements require. However, this new approach created a key trade-off: focusing on individual components meant losing the valuable context provided by full theorem statements. Based on this, we identify two main limitations in the current approach to retrieval-augmented autoformalization. First, informal statements are often dense and multifaceted. This underlying complexity of queries makes them suboptimal as direct queries for retrieving the precise, atomic definitions required for formalization. Second, even when the correct formal definitions are retrieved, models often lack the knowledge of contextual usage required to correctly structure and integrate them into the formal statement.
@@
In information retrieval, query enhancement techniques like query expansion (Chan et al., 2024), pseudo-document generation (Gao et al., 2023), and neural query rewriting (Wang et al., 2025) have demonstrated the effectiveness of reformulating queries to provide more semantic information. Query decomposition has proven particularly useful for multi-hop question answering as it matches the granularity of the dense query statements with the indexed documents (Ammann et al., 2025). In addition to retrieving correct premises, providing rich context like exemplar proofs can guide proof generation effectively (He et al., 2024; Thakur et al., 2024; Thompson et al., 2025). Despite advances in adjacent domains, these techniques have not yet been systematically applied to dependency retrieval for autoformalization, which still relies on monolithic queries and provides context-free definitions.
@@
We propose Drift, a novel framework depicted in Figure 1, that enhances retrieval-augmented autoformalization by adapting query decomposition and context augmentation to address the unique challenges of theorem autoformalization. To tackle the complexity of informal statements, we first adapt query decomposition to break down statements into a series of simpler, atomic sub-queries. Each sub-query targets a single mathematical concept, transforming a multifaceted retrieval problem into focused, precise searches. To address the challenge of correctly applying the dependencies, we contextualize the retrieved definitions with illustrative theorems, giving the model concrete examples of syntax and application patterns.
@@
Contributions. 1) We introduce Drift 111The code and models are available at https://github.com/huawei-noah/DRIFT. (Decompose, Retrieve, Illustrate, then Formalize Theorems), a decomposition-driven retrieval-augmented formalization framework that breaks down informal mathematical statements into atomic sub-queries and contextualizes retrieved premises with demonstrative theorems, transforming monolithic retrieval into a process aligned with the dependency structure of formal mathematics. 2) Our experiments establish new state-of-the-art in dependency retrieval and autoformalization on ProofNet and ConNF across models, with exceptional performance on the out-of-distribution ConNF benchmark. 3) Through systematic analysis, we establish that the utility of retrieved dependencies is conditioned on the gap between a model’s parametric knowledge and the statement’s complexity. These insights reveal critical design considerations for retrieval-augmented systems and point toward the necessity of adaptive strategies that can assess when external knowledge genuinely complements model capabilities.
@@ 2 Related Work
Retrieval-augmented Autoformalization. Early retrieval-augmented autoformalization methods retrieved similar theorems as few-shot examples. For instance, ProofNet (Azerbayev et al., 2023) employs k-NN search, while MS-RAG (Zhang et al., 2024) uses informal-to-informal retrieval with iterative refinement. LTRAG (Hu et al., 2025) retrieves thought-guided theorem pairs for neuro-symbolic formalization. In a key development, Liu et al. (2025) established the “dependency retrieval” paradigm, a premise selection task specialized for autoformalization: given an informal statement, retrieve the precise set of formal objects and definitions that are required for its autoformalization from a library (e.g., Mathlib). While RAutoformalizer (RAuto) (Liu et al., 2025) demonstrated improvements over non-retrieval methods, evaluation revealed a significant gap compared to oracle systems with ground-truth dependencies. CRAMF (Lu et al., 2025) attempts conceptual mapping between abstraction levels. Critically, all existing approaches treat complex statements as monolithic queries, failing to identify distinct mathematical concepts within them.
Query Decomposition and Enhancement. Query enhancement has proven effective across retrieval tasks. Query2Doc (Wang et al., 2023) and HyDE (Gao et al., 2023) generate pseudo-documents to expand semantic coverage. SemanticSearch (Gao et al., 2024) augments the informal statement by prompting an LLM to translate the query into a detailed statement with both informal and formal expressions. However, they only evaluated with similar theorem retrieval but not on the downstream formalization. More relevant to our work, query decomposition has shown success in multi-hop question answering (Ammann et al., 2025), where breaking complex questions into sub-queries improves retrieval of distinct information aspects. Zhao et al. (2023) and Jiang et al. (2022) applied similar decompositions to theorem proving, showing the effectiveness of divide and conquer in formal math. Despite these successes, no prior work has applied decomposition to informal mathematical statements for autoformalization.
@@ 3 Methodology
We introduce Drift (Decompose, Retrieve, Illustrate, then Formalize Theorems), a novel four-stage method designed to address the two main limitations of previous retrieval-augmented formalization methods: 1) the complexity of informal statements and 2) the lack of demonstrative examples for retrieved formal objects. As a first step, an LLM decomposes the informal statement into a set of smaller, granular sub-queries (§3.1). These queries then guide the retrieval of dependent premises from a formal library such as Mathlib (§3.2). In a subsequent, bottom-up illustration step, we find theorems that utilize the retrieved premises, providing in-context demonstrations of their application (§3.3). Finally, the LLM formalizes the original statement, conditioned on all retrieved premises and theorems (§3.4). This process is visualized in Figure 1.
@@ 3.1 Decompose
Standard retrieval methods often treat complex informal statements as monolithic queries and simply embed the entire statement. This approach disregards the rich semantic structure within a statement, which may contain multiple distinct mathematical concepts. Their complexity means that a statement could be under-specified, ambiguous, and/or simply too information-dense to be used as is. Compressing the entire statement’s meaning into a single dense vector creates a representational bottleneck, risking the loss of nuanced details and focusing the retrieval on only the most salient concepts. We hypothesize that by decomposing an informal statement into its constituent concepts, we can perform a more granular and accurate retrieval for each concept.
@@
where , the number of sub-queries, varies for each informal statement and is determined dynamically by the decomposer. Each sub-query is designed to isolate a single mathematical concept. As illustrated in Figure 1, the component is a natural language phrase describing the concept (e.g., “A prime number of the form is a prime that leaves remainder 3 when divided by 4”) while is a predicted formal representation for that concept (e.g., “ : with the conditions Nat.Prime and % , where % denotes the modulo operation on natural numbers.”). Appending this predicted formal name serves a dual purpose. First, it probes the LLM’s parametric knowledge, providing a syntactic “anchor” for the concept. Second, it allows the retriever to jointly leverage the semantics of the natural language phrase and the syntactic cues from to identify the correct premise even in the presence of minor inaccuracies in the predicted formal representations.
@@ 3.2 Retrieve
The Retrieve module is designed to identify dependent premises from the library that correspond to the concepts isolated in each sub-query. This one-to-one mapping is visualized in the panel ② of Figure 1, which shows each sub-query () being linked to a formal object (e.g., PNat.Prime). To accomplish this, we implement a dense passage retriever (Karpukhin et al., 2020) using a BGE-M3 (Chen et al., 2024) encoder (), finetuned on the dependency retrieval task as introduced by Liu et al. (2025). This training objective aligns the informal statements with their formal dependencies, thereby making semantic similarity a strong proxy for logical dependency. We retrieve dependencies by encoding queries and formal library objects into a shared -dimensional embedding space. The vector representations for all formal objects are pre-computed in an offline step and stored in an efficient search index. At inference, each sub-query is encoded into a vector and the closest dependent premise is identified by finding the library object that maximizes the cosine similarity score, , with the query vector. This is formally defined as:
@@ 3.3 Illustrate
Retrieving useful formal definitions is a necessary first step; however, it is not sufficient for successful autoformalization. For instance, a retrieved definition like “def ZMod : → Type” tells the model what the concept is, but provides no further guidance on its practical application, such as the syntax for declaring a variable of that type (a : ZMod p). This gap between definition and usage is a primary source of syntactic and structural errors in LLM-generated formal statements.
The Illustrate module is designed to bridge this gap by providing examples of formal object usage, visualized in Figure 1 (panel ③). Given a premise like ZMod from the “Retrieve” step, the module selects illustrative theorems that demonstrate the correct application of ZMod, such as “Theorem 2”. The module takes the set of retrieved premises , and a budget , as input, and outputs a small set of illustrative theorems , where . The selection process is a greedy algorithm designed to maximize the coverage of the input premises as follows. First, we compile a candidate set of all theorems in the library that utilize at least one of the retrieved premises .
In order to ensure relevance and provide a deterministic tie-breaker, we pre-sort this candidate set in descending order of semantic similarity . For each theorem , this similarity score is the cosine similarity between its informal statement and the original informal statement , computed using the same encoder from the retrieval stage: .
@@ 4.1 Benchmarks
We evaluate Drift on three distinct autoformalization benchmarks to test its in-distribution, self-contained, and out-of-distribution performance. While the experiments are conducted in Lean, our framework is language-agnostic and adaptable to other formal systems with structured libraries.
@@
MiniF2F (Self-Contained). We use the MiniF2F-test set (Zheng et al., 2021) to evaluate the framework on self-contained problems. This benchmark consists of 224 Olympiad-style theorems with a notably low average of just 0.43 dependencies from Mathlib.222We removed 20 examples from the dataset that were either duplicated or failed to compile (Lean v4.18.0). MiniF2F-test serves as a boundary condition, testing the model’s core formalization capabilities and its robustness against potentially distracting context when retrieval is not strictly necessary.
ConNF (Out-of-Distribution). The OOD challenge refers to evaluation scenarios where neither the retrieval model nor the formalization model has been exposed to the formal objects used in the test data.333The knowledge cutoff dates of the models we used are June 2024, March 2025, and July 2025, for GPT-4.1, Claude-Opus-4, and DeepSeek-V3.1, respectively. From our zero-shot results on ConNF, we hypothesize that the models have not been extensively trained on this benchmark, though there is a risk of exposure to the formalizer which cannot be controlled or confirmed. We therefore evaluate on ConNF, a benchmark curated by Liu et al. (2025) through topological informalization to test OOD generalization. This benchmark is based on the con-nf library and is not integrated with Mathlib.444The con-nf library is available at https://github.com/leanprover-community/con-nf. It formalizes a consistency proof for Quine’s New Foundations (Quine, 1951), established by Holmes & Wilshaw (2015) and contains 961 research-level theorems requiring an average of 3.92 premises from its distinct library of 1,348 formal objects. ConNF rigorously tests Drift’s generalization to a novel mathematical domain (Zhang et al., 2024).
@@ 4.2 Baselines
We evaluate our proposed framework against three key baselines representing zero-shot, state-of-the-art retrieval, and oracle-level performance. We note that concurrent work, CRAMF (Lu et al., 2025), was not publicly available for comparison at the time of our experiments. The no retrieval (zero-shot) baseline establishes a performance floor. The autoformalization model receives only the informal statement as input, without access to any retrieved context. DPR (RAuto) represents the current state-of-the-art. We compare our method to the dependency retrieval module from the RAutoformalizer framework (Liu et al., 2025), which is a finetuned dense passage retriever (DPR) (Karpukhin et al., 2020).555Model available at https://huggingface.co/purewhite42/dependency_retriever_f. The top-5 retrieved premises are provided to the formalizer model as augmented context. The oracle* (retrieval) setting provides an approximate upper bound for retrieval. The model is provided with ground-truth dependencies () for each problem, simulating a perfect retriever, as defined by Liu et al. (2025). We mark this setting with an asterisk () to denote that this oracle is, in fact, imperfect. This is because the provided dependencies are not necessarily optimal or exhaustive for formalization and do not necessarily lead to the best autoformalization performance. As we discuss in Section 5, some of our settings actually outperform this imperfect oracle.